OpenStudy (smarty):

Solve for X 3x + 2y + z = 6 x + y + 3z = –5 4x + y – z = 10

OpenStudy (anonymous):

\[(3x + 2y + z) + (4x + y - z) = 16 => 7x + 3y = 16 \] Multiply the third equation by 3, then add to the second, we'll get a new equation. Using it and what we got above to solve for x and y. Then find z.

OpenStudy (smarty):

12x+3y-3z=30

OpenStudy (smarty):

@cwrw238

OpenStudy (cwrw238):

x + y + 3z = –5 12x+3y-3z=30 add these 2 equations 13x + 4y = 25 7x + 3y = 16 - the equation from tvcong

OpenStudy (cwrw238):

now multiply ist equation by -3 and 2nd by 4 -39x - 12y = -75 28x + 12y = 64 adding: -11x - -11 x = 1 plug x = 1 into 7x + 3y = 16: 7 + 3y = 16 3y = 9 y = 3 now plug x = 1 and y = 3 into one of the original equations to find z